OpenStudy (amariiiaaa):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 1, -6>; v = <6, 7, -3>; w = <-7, 5, 2>

OpenStudy (3mar):

May I help?

OpenStudy (amariiiaaa):

Yes please

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Can I upload a file to you?

OpenStudy (amariiiaaa):

I'm still confused

OpenStudy (3mar):

Simply man! \[\large v = (6, 7, -3) ~~ and ~~w =(-7, 5, 2)\] \[\large v.w=(6, 7, -3).(-7, 5, 2)=(6*-7)+(7*5)+(-3*2)\] \[\large v.w=-42+35-6=\Huge-13\]

OpenStudy (amariiiaaa):

Thank you

OpenStudy (3mar):

Satisfied?

OpenStudy (3mar):

Thank you for learning!

OpenStudy (3mar):

Thank you for the medal!